OpenStudy (anonymous):

2. Find the slope of the line described by: x - 7*y = -6 3. Tell whether the equation: -x + 5*y = -6 represents a direct variation. Explain why/why not. 4. Find the x- and y-intercepts of -x + 2y = 20

OpenStudy (solomonzelman):

2. x - 7y = -6 x - 7y - x = -6 -x -7y= -6 -x (-1) 7y = (-1) (6+x) 7y=x+6 y=1/7x+6/7 The Slope is 1/7 3. -x + 5y = -6 -x + 5y +x = -6 + x 5y = x -6 y = 1/5x +6 It is NOT a direct variation, because it has a y-intercept. 4. -x + 2y = 20 -x + 2y + x = 20 + x 2y = x + 20 y = 1/2x +10 find x intercept find y intercept y = 1/2x +10 y = 1/2x +10 (0) = 1/2x +10 y = 1/2(0) +10 0= 1/2x+10 y= 0 + 10 = 10 0 -10 = 1/2x+10 - 10 -10 = 1/2x (both sides times 2) -20 =x x intercept = -20 y intercept =10